# Yixin Liu1 2  Lie Lu2  Jihui Jin2  Lichao Sun1  Andrea Fanelli2

### v2

@@ Abstract
The rapid proliferation of generative audio synthesis and editing technologies has raised significant concerns about copyright infringement, data provenance, and the spread of misinformation through deepfake audio. Watermarking offers a proactive solution by embedding imperceptible, identifiable, and traceable marks into audio content. While recent neural network-based watermarking methods like WavMark and AudioSeal have improved robustness and quality, they struggle to achieve both robust detection and accurate attribution simultaneously. This paper introduces Cross-Attention Robust Audio Watermark (XAttnMark), that bridges this gap by leveraging partial parameter sharing between the generator and the detector, a cross-attention mechanism for efficient message retrieval, and a temporal conditioning module for improved message distribution. Additionally, we propose a psychoacoustic-aligned temporal-frequency masking loss that captures fine-grained auditory masking effects, enhancing watermark imperceptibility. Our approach achieves state-of-the-art performance in both detection and attribution, demonstrating superior robustness against a wide range of audio transformations, including challenging generative editing with strong editing strength. Project webpage is available at https://liuyixin-louis.github.io/xattnmark/.
@@ 1 Introduction
With the rapid development of generative audio synthesis and editing techniques, anyone can now easily edit and re-synthesize audio content OpenAI (2024); Li et al. (2024); Copet et al. (2024). While it democratizes the creative process and enables new applications, it also brings serious concerns for copyrighted data misuse, data provenance and authenticity Pan et al. (2023); Shoaib et al. (2023); Park et al. (2023). A notable example is the recent surge of deepfake audio and video, where actors have been using deepfake techniques to impersonate and create fake speech and video content of online politicians or public figures, with the malicious intent to spread misinformation and manipulate public opinion Verma et al. (2024); Wenger et al. (2021); Buo (2020); Bilika et al. (2023). Furthermore, beyond deepfake threats, the unauthorized exploitation of copyrighted content is also a growing concern in the AI industry Singer (2024); Qiwei et al. (2024); Brigham et al. (2024). These days, many content creators fall victim of copyright infringement due to the unauthorized use of their content for AI training and editing (Office, 2023; Abbott & Rothman, 2023). Original content is now being exploited and modified at scale in a way that makes it hard to track data provenance Cho (2024); Robinson (2024); Vermillio (2024). Among the various solutions to track audio provenance and guarantee artists’ protection Ren et al. (2024); Desai & Riedl (2024); Liu et al. (2024c), watermarking stands out as one of the most effective proactive approaches. It involves embedding imperceptible perturbations into the audio that are both identifiable and traceable. Watermarking enables two key processes: detection, which verifies the presence of the watermark in an audio file, and attribution, which involves decoding a message that uniquely identifies the original creator.
@@
Initialized by WavMark Chen et al. (2023) and the seminal work of AudioSeal San Roman et al. (2024), using end-to-end deep neural networks for learning to watermark audio content has demonstrated stronger robustness with minor quality degradation compared to the state-of-the-art hand-crafted watermarking method Westerfeld (2020). This is especially true under more challenging audio editing transformations, like EnCodec (Defossez et al., 2022). WavMark proposes to use an invertible neural network architecture with a 16-bit synchronization code and a 16-bit message code to jointly conduct detection and attribution. However, the brute-force decoding approach in WavMark is inefficient, and the invertible architecture limits the watermarking capacity under more challenging transformations (Chen et al., 2023). Further work is represented by AudioSeal (San Roman et al., 2024), which utilizes a disjointed generator-detector framework with a detection head and a message decoding head split to improve robustness against advanced transformations. Nevertheless, AudioSeal improves detection at the cost of reduced accuracy in attribution. In a nutshell, whether the neural-network-based watermarking can achieve both robust detection and attribution is still an unsolved puzzle.
@@
This paper identifies the following key factors to bridge the gap: first, motivated by the shared-parameter architecture advantage of WavMark for boosting learning efficiency, and the disjointed generator-detector architecture of AudioSeal for robust capability, we introduce a blended architecture of partial parameter sharing between generator and detector, which jointly achieve both efficient learning and robustness. Specifically, we propose a cross-attention module that leverages a shared embedding table to facilitate message decoding in the detector part. Secondly, we design a simple yet effective conditioning mechanism that distributes the message temporally before injection, which further improves the learning efficiency. With these two key components, we obverse significant boost in both robust detection and attribution performance. Furthermore, to enforce stronger constraints for watermark imperceptibility, we introduce a new per-tile temporal-frequency (TF) masking loss. Specifically, we first compute masking energy with an asymmetric 2D kernel, identify the masked regions and then use the masking energy as a weighting factor for computing a TF-weighted $\ell_{2}$ loss in the mel-spectrogram domain. With these efforts, we demonstrate state-of-the-art robustness across a wide range of audio editing transformations, while preserving superior perceptual quality (see Figure 1). Furthermore, under the more challenging task of generative model editing, we demonstrate that XAttnMark is the only watermarking approach that can conduct watermark detection even when edits of strong strength are applied. We summarize our main contributions as follows:
@@
  • Blending the architecture advantages of previous works, we introduce partial parameter sharing between the neural generator and the detector, with an embedding table as the bridge and a cross-attention module in the detector as the core, to allow for more efficient learning and accurate message retrieval. Furthermore, we introduce a simple yet effective message conditioning module that distributes the latent message temporally, boosting the attribution learning efficiency.
@@
  • To increase the watermarking quality, we introduce a new psychoacoustic-inspired temporal-frequency (TF) masking loss that captures per-tile masking effects. We compute masking energy with an asymmetric 2D kernel, identify the masked TF tiles, and downweight these tiles with a TF-weighted $\ell_{2}$ loss, achieving more imperceptible watermarking.
@@
  • We empirically show that our approach can achieve state-of-the-art performance in both detection and attribution with comparable perceptual quality and superior robustness. Furthermore, testing in a zero-shot manner on unseen generative editing transformations, we show that our approach is the only one that can still successfully conduct watermark detection even with strong editing strength.
@@ 2 Related Work
Audio Watermarking. Audio watermarking has evolved significantly from traditional signal processing to modern deep learning approaches. Early rule-based methods focused on embedding watermarks in time or frequency domains through hand-crafted techniques Zhang (2020); Hu et al. (2020); Zhang & Han (2017); Qin et al. (2022). A notable example is AudiowMark Westerfeld (2020), which embeds a 128-bit message using convolutional coding and selective frequency band modifications. While sophisticated, these hand-crafted approaches often struggle with robustness against challenging transformations like neural audio codecs (Defossez et al., 2022). Deep neural networks (DNNs) have enabled more robust end-to-end watermarking systems that can generalize to unseen transformations (San Roman et al., 2024; Chen et al., 2023; Liu et al., 2023a). WavMark Chen et al. (2023) introduced an invertible neural architecture for joint detection and attribution with 16-bit synchronization codes. While achieving strong performance, its brute-force decoding and architectural constraints limit scalability. AudioSeal San Roman et al. (2024) addressed these limitations with a generator-detector design with separate detection and message decoding. However, this improved detection came at the cost of attribution accuracy. Our work focuses on enabling both robust detection and accurate attribution through a more efficient architecture design with psychoacoustic-inspired quality loss.
@@
Source Attribution. A central objective in copyright protection is the ability to trace and verify the origin of creative works, which remains challenging in the realm of generative audio. Recent efforts have highlighted the necessity of robust source attribution mechanisms that work reliably across different transformations. For instance, Agnew et al. Agnew et al. (2024) performed an extensive audit of popular audio datasets and revealed serious intellectual property infringements, underscoring the urgency for transparent dataset documentation and reliable authorship checks. In the music domain specifically, Barnett et al. Barnett et al. (2024) advanced source attribution by leveraging audio embeddings to identify influential training data in generative music models, enabling a more transparent “musical roots” analysis. Such embedding-based similarity checks align with the broader push for dataset auditing, as reflected in Du et al. Du et al. (2024), who argue for holistic copyright auditing mechanisms throughout the existing machine learning processes. In this work, we propose a neural watermarking system that achieves SoTA message decoding performance, which is an essential step toward robust source attribution.
@@
Figure 2: System Diagram for XAttnMark. XAttnMark consists of a watermark generator and a watermark detector, with a shared embedding table that facilitates message decoding through a cross-attention module. In the generator part, we first employ an encoder network to encode the audio latent and then apply a temporal modulation to hide the message. The modulated latent is then fed into a decoder to produce the watermark residual. In the detector part, a linear detection head is used for detecting the presence of watermarks, and a cross-attention module with the shared embedding table is used for message decoding.
@@ 3 Preliminaries
Audio Watermarking. Audio watermarking systems typically comprise two primary components: a generator $\mathcal{G}$ to embed watermark, and a detector $\mathcal{D}$ for recovering them. Let $\bm{x}\in\mathbb{R}^{T}$ be an audio signal of length $T$, and let $\bm{w}\in\{0,1\}^{K}$ be a binary watermark sequence. The generator $\mathcal{G}:\mathbb{R}^{T}\times\{0,1\}^{K}\rightarrow\mathbb{R}^{T}$ outputs a watermarked signal $\bm{x}_{w}=\mathcal{G}(\bm{x},\bm{w})$, which should ideally preserve the audio’s perceptual quality. The detector $\mathcal{D}$ is then responsible for two tasks. First, its detection head $\mathcal{D}^{\mathrm{det}}:\mathbb{R}^{T}\rightarrow[0,1]$ produces a probability indicating whether an input contains a valid watermark, which yields a final decision of presence or absence with a threshold. Second, its decoder head $\mathcal{D}^{\mathrm{msg}}:\mathbb{R}^{T}\rightarrow[0,1]^{K}$ returns a vector indicating the probability of each bit being 1, which can be thresholded to reconstruct the embedded message. In practical settings, the watermarked audio may undergo various transformations $\mathcal{T}(\cdot)$ such as compression and cropping, resulting in a distorted signal $\bm{x}_{w}^{T}=\mathcal{T}(\bm{x}_{w})$. Across different transformations, an ideal detector should be able to robustly detect and decode watermarks. In essence, the detection head should output probabilities close to 1 for watermarked signals (and close to 0 otherwise), while the decoder head should recover a bit sequence $\tilde{\bm{w}}$ matching $\bm{w}$. In this paper, we focus on the per-sample level detection and attribution, which is a common and practical setting in real-world watermarking applications Liu et al. (2024a).
@@
where $\bm{x}_{w}=\mathcal{G}(\bm{x},\bm{w})$ and $\bm{x}_{w}^{T}=\mathcal{T}(\bm{x}_{w})$. During the inference time, the threshold $\tau$ is usually tuned on the validation set to control a certain level of false positive rate.
Meanwhile, the decoder head produces $K$ probabilities $\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})=[\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})_{1},\dots,\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})_{K}]$, where each entry indicates the likelihood of a particular bit being 1. Its training objective is to maximize the probability of correctly predicting each bit of $\bm{w}$, which can be expressed as:
@@
By thresholding each $\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})_{k}$ at around 0.5, the decoded bit string $\hat{\bm{w}}$ is obtained, with the learning objective seeking $\hat{\bm{w}}=\mathbf{w}$ despite the distortion layer $\mathcal{T}$ over $\mathbf{x}_{w}$.
@@
To contextualize our architectural improvements, we first analyze AudioSeal’s framework, which pioneered the disjointed generator-detector paradigm for neural watermarking. While it achieves strong detection robustness, its attribution limitations motivate key aspects of XAttnMark.
@@
Disjointed Generator–Detector Architecture. AudioSeal employs two separate networks for watermark generation and detection: given audio $\bm{x}$, the generator $\mathcal{G}=\{\mathcal{E}_{\text{gen}},\mathcal{J}_{\text{gen}}\}$ is composed of an encoder $\mathcal{E}_{\text{gen}}$ that encodes the audio into compact latent $\bm{h_{x}}$, and a decoder $\mathcal{J}_{\text{gen}}$ that decodes the latent into the watermarked perturbation. The audio latent is obtained via $\bm{h_{x}}=\mathcal{E}_{\text{gen}}(\bm{x})\in\mathbb{R}^{t^{\prime}\times H}$, where $t^{\prime}=\lfloor T/\alpha\rfloor$, $\alpha$ is the temporal downsampling factor, and $H$ is the latent dimension. The secret message $w$ is injected into the latent space with a message encoder $\mathcal{M}$ that maps $\bm{w}$ to its latent $\bm{h}_{w}=\mathcal{M}(\bm{w})$. The modulated latent waveform is obtained via direct addition $\bm{h}(\bm{x},\bm{w})=\bm{h}_{x}+\bm{h}_{w}$, following with the decoder network $\mathcal{J}_{\text{gen}}$ to produce the predicted watermarked perturbation $\bm{\delta}_{\bm{w}}=\mathcal{J}_{\text{gen}}(\bm{h}(\bm{x},\bm{w}))\in\mathbb{R}^{T\times 1}$, which is later applied on the original audio to produce the watermarked audio $\bm{x}_{w}=\bm{x}+\bm{\delta}_{\bm{w}}$. Note that the decoder model $\mathcal{J}_{\text{gen}}$ shares a symmetric structure with the encoder $\mathcal{E}_{\text{gen}}$, with layers of residual transposed convolution for temporal upsampling. The watermark detector $\mathcal{D}=\{\mathcal{E}_{\text{det}},\mathcal{O}^{\mathrm{dec}}\}$ is composed of an encoder $\mathcal{E}_{\text{det}}$ (sharing similar structure with $\mathcal{E}_{\text{gen}}$ but with an additional padding to match original length $T$) and a unified decoding head $\mathcal{O}^{\mathrm{dec}}\in\mathbb{R}^{H\times(1+K)}$ for both detection and message decoding. Given potential watermarked audio $\bm{x}_{w}^{T}$, the encoder $\mathcal{E}_{\text{det}}$ first processes the audio to obtain the latent $\bm{{\tilde{h}}}_{x}=\mathcal{E}_{\text{det}}(\bm{x}_{w}^{T})\in\mathbb{R}^{T\times H}$. Then the decoding head $\mathcal{O}^{\mathrm{dec}}$ is applied on the latent to obtain the logits for both detection and message decoding, i.e., $\mathcal{O}^{\mathrm{dec}}(\bm{{\tilde{h}}}_{x})\in\mathbb{R}^{T\times(1+K)}$.
@@
Multi-bit Message Conditioning $\mathcal{M}$. Similar to the common practice of using learned embeddings for conditional generation (Peebles & Xie, 2023), AudioSeal uses an embedding table of only $2K$ entries, $\mathbf{E}\in\mathbb{R}^{2K\times H}$, to encode the total $2^{K}$ possible message states, which significantly reduces the storage complexity. Specifically, each bit $w_{j}\in\{0,1\}$ maps to index $I_{j}=2j+w_{j}$, allowing separate representations for 0/1 states in each bit position with distinct embedding vectors. Given the position sequence $\bm{p}=[I_{1},\dots,I_{K}]$, the retrieved $K$ embedding vector sequence is $\mathbf{V}(\bm{w})=[\mathbf{E}_{I_{1}},\dots,\mathbf{E}_{I_{K}}]\in\mathbb{R}^{K\times H}$. Then an average pooling operation on $\mathbf{V}(\bm{w})$ is applied to obtain the final message latent $\bm{h}_{w}=\frac{1}{K}\sum_{j=1}^{K}\mathbf{E}_{I_{j}}\in\mathbb{R}^{H}$. The message latent is repeated along the temporal axis and then added to the waveform latent, leading $\bm{h}(\bm{x},\bm{w})=\bm{h}_{x}+\bm{h}_{w}$.

### v3

# XAttnMark: Learning Robust Audio Watermarking with Cross-Attention

@@ Abstract
The rapid proliferation of generative audio synthesis and editing technologies has raised serious concerns about copyright infringement, data provenance, and the spread of misinformation via deepfake audio. Watermarking offers a proactive solution by embedding imperceptible yet identifiable and traceable signals into audio content. While recent neural network-based watermarking methods like WavMark and AudioSeal have improved robustness and quality, they struggle to jointly optimize both robust detection and accurate attribution. This paper introduces Cross-Attention Robust Audio Watermark (XAttnMark), which bridges this gap by leveraging partial parameter sharing between the generator and the detector, a cross-attention mechanism for efficient message retrieval, and a temporal conditioning module for improved message distribution. Additionally, we propose a psychoacoustic-aligned time-frequency (TF) masking loss that captures fine-grained auditory masking effects, improving watermark imperceptibility. XAttnMark achieves state-of-the-art performance in both detection and attribution, demonstrating superior robustness against a wide range of audio transformations, including challenging generative editing at varying strengths. This work advances audio watermarking for protecting intellectual property and ensuring authenticity in the era of generative AI.
Audio Watermarking, Deep Learning, Attribution, Robustness
@@ 1 Introduction
With the rapid development of generative audio synthesis and editing techniques, anyone can now easily edit and synthesize audio content (OpenAI, 2024; Li et al., 2024; Copet et al., 2024; Cao et al., 2025; Zhou et al., 2024). While it democratizes the creative process and enables new applications, it also raises serious concerns regarding unauthorized use of copyrighted data, data provenance, and authenticity (Pan et al., 2023; Shoaib et al., 2023; Park et al., 2023; Liu et al., 2024c; Meerza et al., 2025; Liu et al., 2024e). A notable example is the recent surge in deepfake audio and video, where malicious actors use generative techniques to impersonate and create fake speech and video content of online politicians or public figures, to spread misinformation and manipulate public opinion (Verma et al., 2024; Wenger et al., 2021; Buo, 2020; Bilika et al., 2023). Furthermore, beyond deepfake threats, the unauthorized exploitation of copyrighted content is also a growing concern in the AI industry (Singer, 2024; Qiwei et al., 2024; Brigham et al., 2024; Cui et al., 2024; Guan et al., 2024; Liu et al., 2024d). Nowadays many content creators are subject to copyright infringement due to the unauthorized use of their content for AI training and editing (Office, 2023; Abbott and Rothman, 2023). Original content is increasingly exploited and modified at scale, making it difficult to track data provenance (Cho, 2024; Robinson, 2024; Vermillio, 2024). As highlighted in Content Authenticity Initiative (CAI) (Wikipedia contributors, 2025), the challenge of combating disinformation and verifying content authenticity in the digital age requires robust standards for provenance metadata. Among the various solutions to track audio provenance and guarantee artists’ protection (Ren et al., 2024; Desai and Riedl, 2024), watermarking stands out as one of the most effective proactive approaches. It involves embedding imperceptible perturbations into the audio that are both identifiable and traceable. Watermarking enables two key processes: detection, which verifies the presence of the watermark in an audio file, and attribution, which involves decoding a message that uniquely identifies the original creator.
@@
Initialized by WavMark (Chen et al., 2023) and the seminal work of AudioSeal (San Roman et al., 2024), using end-to-end deep neural networks for learning to watermark audio content has demonstrated stronger robustness with minor quality degradation compared to the state-of-the-art hand-crafted watermarking method (Westerfeld, 2020). This is evident under challenging edits, such as EnCodec (Défossez et al., 2023). WavMark proposes an invertible neural network architecture with a 16-bit synchronization code and a 16-bit message code to jointly conduct detection and attribution. However, the brute-force decoding approach in WavMark is inefficient, and the invertible architecture limits the watermarking capacity under more challenging transformations (Chen et al., 2023). AudioSeal (San Roman et al., 2024) addresses these issues by decoupling the generator and detector and adding separate heads for detection and message decoding, which improves robustness to advanced transformations. However, this design boosts detection at the cost of lower attribution accuracy. In summary, whether neural-network-based watermarking can achieve both robust detection and attribution is still an open problem.
@@
This paper identifies two key factors to bridge this gap: first, motivated by the shared-parameter architecture advantage of WavMark for boosting learning efficiency, and the disjoint generator-detector architecture of AudioSeal for robust capability, we introduce a blended architecture of partial parameter sharing between generator and detector, which jointly achieves both efficient learning and robustness. Specifically, we propose a cross-attention module that leverages a shared embedding table to facilitate message decoding in the detector part. Second, we design a simple yet effective conditioning mechanism that distributes the message temporally before injection, which further improves the learning efficiency. With these two key components, we observe significant gains in both detection robustness and attribution accuracy. To further improve watermark imperceptibility, we introduce a new per-tile TF masking loss. Specifically, we first compute masking energy with an asymmetric 2D kernel, identify the masked regions, and then use the masking energy as a weighting factor for computing a TF-weighted $\ell_{2}$ loss in the mel-spectrogram domain. With these efforts, we demonstrate state-of-the-art robustness across a wide range of audio editing transformations, while preserving superior perceptual quality (see Figure 1). Furthermore, under the more challenging task of generative model editing, we demonstrate that our approach is the only watermarking approach among the evaluated methods that can conduct watermark detection even when edits of strong strength are applied. We summarize our main contributions as follows:
@@
  • Combining architectural advantages from prior works, we design partial parameter sharing between the neural generator and the detector, with an embedding table as the bridge and a cross-attention module in the detector as the core, to allow for more efficient learning and accurate message retrieval. Furthermore, we introduce a simple yet effective message conditioning module that distributes the latent message temporally, boosting the attribution learning efficiency.
@@
  • To improve perceptual quality, we introduce a new psychoacoustic-inspired time-frequency masking loss that captures per-tile masking effects. We compute masking energy with an asymmetric 2D kernel, identify the masked TF tiles, and assign lower loss weights to those tiles with a TF-weighted $\ell_{2}$ loss, achieving more imperceptible watermarking.
@@
  • We empirically show that our approach can achieve state-of-the-art performance in both detection and attribution with comparable perceptual quality and superior robustness. Furthermore, testing in a zero-shot manner on unseen generative editing transformations, our approach is the only one among the evaluated methods that maintains non-trivial detection under strong generative edits.
@@ 2 Related Work
Audio Watermarking. Audio watermarking has evolved significantly from traditional signal processing to modern deep learning approaches. Early rule-based methods focused on embedding watermarks in time or frequency domains through hand-crafted techniques (Zhang, 2020; Hu et al., 2020; Zhang et al., 2019; Qin et al., 2024). A notable example is AudiowMark (Westerfeld, 2020), which embeds a 128-bit message using convolutional coding and selective frequency band modifications. Although carefully engineered, hand-crafted methods often degrade under neural codec compression (Défossez et al., 2023). Deep neural networks (DNNs) have enabled more robust end-to-end watermarking systems that can generalize to unseen transformations (San Roman et al., 2024; Chen et al., 2023; Liu et al., 2023a). WavMark (Chen et al., 2023) introduced an invertible neural architecture for joint detection and attribution with 16-bit synchronization codes. While achieving strong performance, its brute-force decoding and architectural constraints limit scalability. AudioSeal (San Roman et al., 2024) addressed these limitations with a generator-detector design with separate detection and message decoding. The decoupled design improves detection but reduces attribution accuracy. Our work focuses on enabling both robust detection and accurate attribution through a more efficient architecture design with a psychoacoustic-inspired quality loss.
@@
Source Attribution. A central objective in copyright protection is the ability to trace and verify the origin of creative works, which remains challenging, especially for generative audio systems. Recent efforts have highlighted the necessity of robust source attribution mechanisms that work reliably across different transformations. For instance, Agnew et al. (Agnew et al., 2024) performed an extensive audit of popular audio datasets and revealed serious intellectual property infringements, underscoring the urgency for transparent dataset documentation and reliable authorship checks. In the music domain specifically, Barnett et al. (Barnett et al., 2024) advanced source attribution by leveraging audio embeddings to identify influential training data in generative music models, enabling a more transparent “musical roots” analysis. Such embedding-based similarity checks align with the broader push for dataset auditing, as reflected in Du et al. (Du et al., 2024), who argue for holistic copyright auditing mechanisms throughout the existing machine learning processes. In this work, we propose a neural watermarking system that advances in message decoding performance, marking an essential step toward robust source attribution.
@@
Figure 2: System Diagram for XAttnMark. XAttnMark consists of a watermark generator and a watermark detector, with a shared embedding table that facilitates message decoding through a cross-attention module. In the generator part, we first employ an encoder network to encode the audio latent and then apply a temporal modulation to hide the message. The modulated latent is then fed into a decoder to produce the watermark residual. In the detector part, a linear detection head is used for detecting the presence of watermarks, and a cross-attention module with a shared embedding table is used for message decoding.
@@ 3 Preliminaries
Audio Watermarking. Audio watermarking systems typically comprise two primary components: a generator $\mathcal{G}$ to embed watermark, and a detector $\mathcal{D}$ for recovering it. Let $\bm{x}\in\mathbb{R}^{T}$ be an audio signal of length $T$, and let $\bm{w}\in\{0,1\}^{K}$ be a binary watermark sequence. The generator $\mathcal{G}:\mathbb{R}^{T}\times\{0,1\}^{K}\rightarrow\mathbb{R}^{T}$ outputs a watermarked signal $\bm{x}_{w}=\mathcal{G}(\bm{x},\bm{w})$, which should ideally preserve the audio’s perceptual quality. The detector $\mathcal{D}$ is then responsible for two tasks. First, its detection head $\mathcal{D}^{\mathrm{det}}:\mathbb{R}^{T}\rightarrow[0,1]$ produces a probability indicating whether an input contains a valid watermark, which yields a final decision of presence or absence with a threshold. Second, its message decoding head $\mathcal{D}^{\mathrm{msg}}:\mathbb{R}^{T}\rightarrow[0,1]^{K}$ returns a vector indicating the probability of each bit being 1, which can be thresholded to reconstruct the embedded message. In practical settings, the watermarked audio may undergo various transformations $\mathcal{T}(\cdot)$ such as compression and cropping, resulting in a distorted signal $\bm{x}_{w}^{T}=\mathcal{T}(\bm{x}_{w})$. Across different transformations, an ideal detector should be able to robustly detect and decode watermarks. In essence, the detection head should output probabilities close to 1 for watermarked signals (and close to 0 otherwise), while the message decoding head should recover a bit sequence $\tilde{\bm{w}}$ matching the original watermark message $\bm{w}$. In this paper, we focus on the per-sample level detection and attribution, which is a practical and widely adopted setting in real-world watermarking systems (Liu et al., 2024b).
@@
where $\bm{x}_{w}=\mathcal{G}(\bm{x},\bm{w})$ and $\bm{x}_{w}^{T}=\mathcal{T}(\bm{x}_{w})$. During inference, the threshold $\tau$ is usually tuned on the validation set to control a certain level of false positive rate.
Meanwhile, the message decoding head produces $K$ probabilities $\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})=[\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})_{1},\dots,\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})_{K}]$, where each entry represents the probability that the corresponding bit is 1. Its training objective is to maximize the probability of correctly predicting each bit of $\bm{w}$, which can be expressed as:
@@
By thresholding each $\mathcal{D}^{\mathrm{msg}}(\bm{x}_{w}^{T})_{k}$ at 0.5, the decoded bit string $\hat{\bm{w}}$ is obtained, with the learning objective seeking $\hat{\bm{w}}=\bm{w}$ despite the distortion layer $\mathcal{T}$ applied to $\bm{x}_{w}$.
@@
To contextualize our architectural improvements, we first analyze AudioSeal’s framework, which pioneered the disjoint generator-detector paradigm for neural watermarking. While it achieves strong detection robustness, its attribution limitations motivate key aspects of XAttnMark.
@@
Disjoint Generator–Detector Architecture. AudioSeal employs two separate networks for watermark generation and detection: given audio $\bm{x}$, the generator $\mathcal{G}=\{\mathcal{E}_{\text{gen}},\mathcal{J}_{\text{gen}}\}$ is composed of an encoder $\mathcal{E}_{\text{gen}}$ that encodes the audio into compact latent $\bm{h_{x}}$, and a decoder $\mathcal{J}_{\text{gen}}$ that decodes the latent into the watermarked perturbation. The audio latent is obtained via $\bm{h_{x}}=\mathcal{E}_{\text{gen}}(\bm{x})\in\mathbb{R}^{t^{\prime}\times H}$, where $t^{\prime}=\lfloor T/\alpha\rfloor$, $\alpha$ is the temporal downsampling factor, and $H$ is the latent dimension. The secret message $w$ is injected into the latent space with a message encoder $\mathcal{M}$ that maps $\bm{w}$ to its latent $\bm{h}_{w}=\mathcal{M}(\bm{w})$. The modulated latent waveform is obtained by adding the message latent to the audio latent $\bm{h}(\bm{x},\bm{w})=\bm{h}_{x}+\bm{h}_{w}$, followed by the decoder network $\mathcal{J}_{\text{gen}}$ to produce the predicted watermarked perturbation $\bm{\delta}_{\bm{w}}=\mathcal{J}_{\text{gen}}(\bm{h}(\bm{x},\bm{w}))\in\mathbb{R}^{T\times 1}$, which is later applied on the original audio to produce the watermarked audio $\bm{x}_{w}=\bm{x}+\bm{\delta}_{\bm{w}}$. Note that the decoder model $\mathcal{J}_{\text{gen}}$ shares a symmetric structure with the encoder $\mathcal{E}_{\text{gen}}$, with layers of residual transposed convolution for temporal upsampling. The watermark detector $\mathcal{D}=\{\mathcal{E}_{\text{det}},\mathcal{O}^{\mathrm{dec}}\}$ is composed of an encoder $\mathcal{E}_{\text{det}}$ (sharing a similar structure with $\mathcal{E}_{\text{gen}}$ but with additional padding to match the original length $T$) and a unified decoding head $\mathcal{O}^{\mathrm{dec}}\in\mathbb{R}^{H\times(1+K)}$ for both detection and message decoding. Given potential watermarked audio $\bm{x}_{w}^{T}$, the encoder $\mathcal{E}_{\text{det}}$ first processes the audio to obtain the latent $\bm{{\tilde{h}}}_{x}=\mathcal{E}_{\text{det}}(\bm{x}_{w}^{T})\in\mathbb{R}^{T\times H}$. Then the decoding head $\mathcal{O}^{\mathrm{dec}}$ is applied on the latent to obtain the logits for both detection and message decoding, i.e., $\mathcal{O}^{\mathrm{dec}}(\bm{{\tilde{h}}}_{x})\in\mathbb{R}^{T\times(1+K)}$.
@@
Multi-bit Message Conditioning $\mathcal{M}$. Similar to the common practice of using learned embeddings for conditional generation (Peebles and Xie, 2023), AudioSeal uses an embedding table of only $2K$ entries, $\bm{E}\in\mathbb{R}^{2K\times H}$, to encode the total $2^{K}$ possible message states, which significantly reduces the storage complexity. Specifically, each bit $w_{j}\in\{0,1\}$ for $j\in\{0,\ldots,K-1\}$ maps to index $I_{j}=2j+w_{j}$, allowing separate representations for 0/1 states in each bit position with distinct embedding vectors. Given the position sequence $\bm{p}=[I_{1},\dots,I_{K}]$, the retrieved $K$ embedding vector sequence is $\bm{V}(\bm{w})=[\bm{E}_{I_{1}},\dots,\bm{E}_{I_{K}}]\in\mathbb{R}^{K\times H}$. Then an average pooling operation on $\bm{V}(\bm{w})$ is applied to obtain the final message latent $\bm{h}_{w}=\frac{1}{K}\sum_{j=1}^{K}\bm{E}_{I_{j}}\in\mathbb{R}^{H}$. The message latent is repeated along the temporal axis and then added to the waveform latent, leading $\bm{h}(\bm{x},\bm{w})=\bm{h}_{x}+\bm{h}_{w}$.